OpenStudy (anonymous):

How do I calculate 10 over X equals 20

OpenStudy (anonymous):

\[\frac{ 10 }{ x } = 20\] Cross multiply \[20x = 10 \] now let x be alone \[x = \frac{ 10 }{ 20}\] Simplify by dividing both dinomiator and numerator by 10 \[\frac{ 1 }{ ? }\]

OpenStudy (anonymous):

I meant the final answer is 1\2